OpenStudy (anonymous):

Can you write 1 1/6 as a proper fraction? Pleaseeeeee explain!! :/

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ 1 \frac{1}{6} = 1 + \frac{1}{6} = \frac{6}{6} + \frac{1}{6} = \frac{1 + 6}{6} = \frac{7}{6} \]